# Occlusion-Aware Seamless Segmentation

### v1

@@ 5.4 Ablation Study
To better understand the components of the UnmaskFormer, we conduct ablation studies for the Unmasking Attention (UA)-based backbone, Amodal-oriented Mix (AoMix) method, and Occlusion-aware Fusion (OAFusion) module.
Why UA unmasks occlusions? In Tab. 7, we ablate different patch embedding methods [77, 94] and pooling methods [60]. Compared to the original Patch Embedding (PE) design and the early-stage DPE, we construct the backbone by using a new interleaving arrangement which improves the distortion-aware modeling, yielding the best mAPQ score with a and a gain respectively. Compared to SimPool [60], our occlusion-aware pooling attention better handles object occlusion and obtains higher results in OASS. An example of feature visualization in Fig. 8 shows that our UA can enhance distortion-tolerant predictions of regions where occlusion occurs, highlighted in red color. These results show the effectiveness of the UA backbone in addressing image distortion and object occlusion at the same time.
How AoMix boosts amodal prediction? As shown in Tab. 7, we investigate various approaches to masked image modeling. Compared to using masked images only for the source image (T for S) or the mixed image (T for M), our strategy (AoMix), which utilizes both, achieves the best results. Masking the Thing class regions (AoMix), as opposed to the entire image (W for S&M) or masked by patches (P for S&M), results in improvements with a gain of and in mAPQ respectively. With the carefully devised AoMix, the UnmaskFormer earns significant gains by greatly boosting surrounding parsing seamlessly.
@@
Finally, as shown in Tab. 7, all components of UnmaskFormer cooperate with each other to achieve the best performance.
@@ 5.5 Analysis of Hyperparameters
Analysis of AoMix parameters. In Fig. 10, we conduct experiments to analyze the impact of different scaling parameters in random scaling operation on the model. Excessively large scaling sizes lead to complete occlusion of objects in images, preventing the model from learning original object information. Overly small scaling sizes result in minimal object occlusion, limiting the model’s capacity to make reasonable predictions for occluded regions. Optimal performance is achieved by setting the random scaling range in to , effectively generating diverse masked source images and masked mixed images. This rich variety of amodal-oriented samples bridges the gap between the pinhole and panoramic domains, enhancing the model’s ability to reconstruct invisible regions of occluded objects and boost seamless segmentation.
@@ 6 Conclusion
A Datasets
Dense annotation of BlendPASS. To enhance the accuracy of dense annotations for objects of the Thing class, particularly in representing invisible regions of occluded objects realistically, our labeling process follows the “independent annotation cross-verification voting” workflow. All images are densely annotated by three skilled annotators using the EISeg tool [24] for initial segmentation. Specifically, three annotators densely annotate the full region of each object, and occluded objects are independently annotated by three annotators. Subsequently, cross-verification is conducted among the annotators. In cases where there are slight discrepancies in the annotations of occluded objects, the final annotation is determined through majority voting. For annotations that do not reach a consensus, more iterative annotation processes are employed until an agreement is reached. This annotation workflow aims to ensure the accuracy of dense annotations while maximizing the fidelity of annotations for occluded regions to the true object shape. Annotating panoramas with severe distortion is challenging and extremely time-consuming, requiring approximately 210 minutes per person per image.
@@ B Evaluation Metrics
AP. We follow Pascal VOC [15] and COCO [48] and use average precision, which is computed by averaging the ten equally spaced IoU thresholds from to .
@@
C More Results
C.1 More OASS Results
@@
C.2 More Analysis of Ablation Study
@@
D More Visualization Results
D.1 More Visualization Results of OASS
@@
D.2 More Visualization Results on SynPASS
@@ E Discussion
Future work. The advancements made in Occlusion-Aware Seamless Segmentation (OASS) through our UnmaskFormer open new avenues for future research. In the future, we envision several potential further directions: 1) User-Interactive Image Editing: We propose exploring user interaction in image editing. Future iterations can integrate user-selected areas to separate amodal layers and employ diffusion processes for completion. These interactive features enhance user-friendliness and expand practical applications. 2) Amodal Optical Flow Fusion: Investigate combining our segmentation approach with amodal optical flow techniques [52] to improve temporal consistency in amodal segmentation. This is useful for dealing with temporal occlusions. Also, it is essential for continuous tracking or monitoring applications, potentially revolutionizing dynamic scene processing.
@@
Limitations and potential solutions. While our UnmaskFormer for OASS showcases significant advancements, we acknowledge certain limitations in our current work, alongside potential approaches to address these challenges: 1) Complex Environment Perception: The current model may not handle the complexity of in-the-wild scenes. Continuous improvement is vital for panoramic image processing and scene understanding. We could enhance robustness through diverse data, integration with other sensors like LiDAR or event cameras, and depth-aware transferring learning [8] to enhance occlusion reasoning. 2) Amodal Data Annotation Challenge: Amodal data annotation for panoramic images is challenging. While our dataset covers diverse images captured in cities located on all continents, we could further explore automated amodal annotation for panoramas, reserving manual annotation for data purification and cleaning purposes. This approach can help scale up the availability of annotated data for training and validation, addressing the scarcity of labeled data in this domain.
@@
Societal impacts. In this study, we have introduced a novel task called Occlusion-Aware Seamless Segmentation (OASS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance in the OASS task, primarily due to the intricate nature of occlusion-aware segmentation challenges. To address this, we developed UnmaskFormer, a solution that significantly enhances performance on the OASS benchmark, outperforming existing domain adaptation panoramic and panoptic segmentation methods and achieving promising state-of-the-art results. Nevertheless, given the criticality of dependability in deep learning systems for Advanced Driver-Assistance Systems (ADAS), it is important to note that UnmaskFormer may still encounter misclassifications in challenging occluded regions and biased content, potentially leading to erroneous predictions with adverse societal implications.

### v2

@@ 5.4 Ablation Study
To better understand the components of the UnmaskFormer, we conduct ablation studies on the KITTI360-APSBlendPASS benchmark.
Why UA unmasks occlusions? In Tab. 7, we ablate different patch embedding methods [72, 89] and pooling methods [55]. The AvgPool and SimPool are used to replace the GAP operator in UA. Compared to the original PE design and the early-stage DPE, we construct the backbone by using a new interleaving arrangement which improves the distortion-aware modeling, yielding the best mAPQ score with a and a gain respectively. Compared to SimPool [55], our occlusion-aware pooling attention better handles object occlusion and obtains higher results in OASS. An example of feature visualization in Fig. 8 shows that our UA can enhance distortion-tolerant predictions of regions where occlusion occurs. These results show the effectiveness of the UA backbone in addressing image distortion and object occlusion at the same time.
How AoMix boosts amodal prediction? As shown in Tab. 7, we investigate various approaches to masked image modeling. Compared to using masked images only for the source image (T for S) or the mixed image (T for M), our strategy AoMix, which utilizes both, achieves the best results. Masking the Thing class regions (AoMix), as opposed to the entire image (W for S&M) or masked by patches (P for S&M), results in improvements with a gain of and in mAPQ respectively. With the carefully devised AoMix, the UnmaskFormer earns significant gains by greatly boosting surrounding parsing seamlessly.
@@
Finally, as shown in Tab. 7, all components cooperate with each other to achieve the best performance. More analyses can be found in the supplementary.
@@ 5.5 Analysis of Hyperparameters
We further conduct an analysis of relevant hyperparameters in UnmaskFormer on the KITTI360-APSBlendPASS benchmark.
@@
Analysis of AoMix parameters. In Fig. 10, we conduct experiments to analyze the impact of different scaling parameters in . Excessively large scaling sizes lead to complete occlusion of objects in images, preventing the model from learning original object information. Overly small scaling sizes result in minimal object occlusion, limiting the model’s capacity to make reasonable predictions for occluded regions. Optimal performance is achieved by setting the range in to , effectively generating diverse masked source images and masked mixed images. This variety of amodal-oriented samples bridges the gap between the pinhole and panoramic domains, enhancing the model’s ability to reconstruct invisible regions of occluded objects and boost seamless segmentation.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the Major Research Plan of the National Natural Science Foundation of China under Grant 92148204, the National Key RD Program under Grant 2022YFB4701400, the Hunan Leading Talent of Technological Innovation under Grant 2022RC3063, the Top Ten Technical Research Projects of Hunan Province under Grant 2024GK1010, the Key Research Development Program of Hunan Province under Grant 2022GK2011, and in part by Hangzhou SurImage Technology Company Ltd.
Appendix 0.A Datasets
Dense annotation of BlendPASS. To enhance the accuracy of dense annotations for objects of the Thing class, particularly in representing invisible regions of occluded objects realistically, our labeling process follows the “independent annotation cross-verification voting” workflow. All images are densely annotated by three skilled annotators using the EISeg tool [hao2022eiseg] for initial segmentation. Specifically, three annotators densely annotate the full region of each object, and occluded objects are independently annotated by three annotators. Subsequently, cross-verification is conducted among the annotators. In cases where there are slight discrepancies in the annotations of occluded objects, the final annotation is determined through majority voting. For annotations that do not reach a consensus, more iterative annotation processes are employed until an agreement is reached. This annotation workflow aims to ensure the accuracy of dense annotations while maximizing the fidelity of annotations for occluded regions to the true object shape. Annotating panoramas with severe distortion is challenging and extremely time-consuming, requiring approximately 210 minutes per person per image. Finally, as illustrated in Tab. S1, objects are annotated in the Thing class. We establish a finely labeled dataset, BlendPASS, based on panoramic images containing semantic, instance, and amodal instance labels. All annotations are cross-checked to support five tasks simultaneously: semantic segmentation, instance segmentation, amodal segmentation, panoptic segmentation, and amodal panoptic segmentation.
189 | 6 | 909 | 42 | 18 | 1 | 83 | 38 | 1286 | |
613 | 12 | 842 | 38 | 24 | 2 | 71 | 72 | 1674 | |
| Total | 802 | 18 | 1751 | 80 | 42 | 3 | 154 | 110 | 2960 |
@@ Appendix 0.B Evaluation Metrics
AP. We follow Pascal VOC [everingham2010pascal] and COCO [lin2014microsoft] and use average precision, which is computed by averaging the ten equally spaced IoU thresholds from to .
@@
Appendix 0.C More Results
0.C.1 More Experiment Details
We use the same data augmentation parameters as DACS [67] and set the RCS temperature to to enhance sample frequency for rare classes of the source domain. The EMA decay is set as . The threshold in pseudo-label weight is set as . Moreover, the pseudo-labels in the regions pixels above and pixels below the trained patches are ignored, respectively. Moreover, we adopt the ImageNet feature distance loss from DAFormer with a weight of . For instance and amodal instance branches, we set all loss weights to . During inference, the thresholds for instance and amodal instance are set to . For the retraining of the existing methods [57, 91, 89, 93], we utilize training protocols similar to ours and followed their hyperparameters to ensure fairness. Our experiments are conducted on an NVIDIA Tesla V100 GPU and implemented using PyTorch.
0.C.2 More OASS Results
@@
0.C.3 More Analysis of Ablation Study
@@
Appendix 0.D More Visualization Results
0.D.1 More Visualization Results of OASS
@@
0.D.2 More Visualization Results on SynPASS
@@ Appendix 0.E Discussion
Future work. The advancements made in Occlusion-Aware Seamless Segmentation (OASS) through our UnmaskFormer open new avenues for future research. In the future, we envision several potential further directions: 1) User-Interactive Image Editing: We propose exploring user interaction in image editing. Future iterations can integrate user-selected areas to separate amodal layers and employ diffusion processes for completion. These interactive features enhance user-friendliness and expand practical applications. 2) Amodal Optical Flow Fusion: We intend to further investigate combining our segmentation approach with amodal optical flow techniques [luz2023amodal_optical_flow] to improve temporal consistency in amodal segmentation. This is useful for dealing with temporal occlusions. Also, it is essential for continuous tracking or monitoring applications, potentially revolutionizing dynamic scene processing.
@@
Limitations and potential solutions. While our UnmaskFormer for OASS showcases significant advancements, we acknowledge certain limitations in our current work, alongside potential approaches to address these challenges: 1) Complex Environment Perception: The current model may not handle the complexity of in-the-wild scenes. Continuous improvement is vital for panoramic image processing and scene understanding. We could enhance robustness through diverse data, integration with other sensors like LiDAR or event cameras, and depth-aware transferring learning [chen2023transferring] to enhance occlusion reasoning. Additionally, domain generalization techniques [jiang2023domain] can be employed to improve the model’s adaptability to new and unseen environments, further bolstering its performance in complex scenarios. 2) Amodal Data Annotation Challenge: Amodal data annotation for panoramic images is challenging. While our dataset covers diverse images captured in cities located on all continents, we could further explore automated amodal annotation for panoramas, reserving manual annotation for data purification and cleaning purposes. Moreover, semi-supervised amodal instance segmentation methods [45] can also be leveraged to enhance the efficiency of the annotation process. These approaches can help scale up the availability of annotated data for training and validation, addressing the scarcity of labeled data in the amodal segmentation domain.
@@
Societal impacts. In this study, we have introduced a novel task called Occlusion-Aware Seamless Segmentation (OASS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance in the OASS task, primarily due to the intricate nature of occlusion-aware segmentation challenges. To address this, we have developed UnmaskFormer, a solution that significantly enhances performance on the OASS benchmark, outperforming existing domain adaptation panoramic and panoptic segmentation methods and achieving promising state-of-the-art results. Nevertheless, given the criticality of dependability in deep learning systems for Advanced Driver-Assistance Systems (ADAS), it is important to note that UnmaskFormer may still encounter misclassifications in challenging occluded regions and biased content, potentially leading to erroneous predictions with adverse societal implications.